OpenStudy (anonymous):

Write the equation in the slope-intercept form of the line that has a slope of -3 and contains the point (4,-5)

OpenStudy (anonymous):

slope intercept form is y=mx+b You know the slope, m=-3 and you have x and y Substitute these values into the equation to solve for b.

OpenStudy (anonymous):

-5=-3(4)+b -5 = -12 + b 7=b Now plug the slope, m, and b into the equation y=mx+b

OpenStudy (anonymous):

y=-3(-12)+7?